OpenStudy (anonymous):

Solve each of these equations by substitution. 8x-4y=6 2x-y=9

OpenStudy (anonymous):

8x + 4y = 6 8(5/8) + 4(1/4) = 6 40/8 + 4/4 = 6 5 + 1 = 6 6 = 6

OpenStudy (anonymous):

2x - y = 1 2(5/8) - 1/4 = 1 10/8 - 1/4 = 1 5/4 -1/4 = 1 4/4 = 1 1 = 1